# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 3.1 Experimental Design Framework
We proceed by generalizing the experimental design. Specifically, we allow treatment assignments $W^{i}_{t}$ to vary over time following a Bernoulli distribution with mean $p_{t}$, and define the experimental design as the product distribution $\Pi=p_{0}\times\ldots\times p_{T}$. In this context, we use $\vec{W}_{t}{}:=(W^{1}_{t},\ldots,W^{N}_{t})^{\top}$ to represent the vector containing the treatment assignments for all experimental units at time $t$. This formulation encompasses a broad range of experimental designs, including staggered roll-out design (Xiong et al., 2024a ), micro-randomized trials (Li and Wager, 2022a ), and switchback experiments (Bojinov et al., 2023).
@@
In Appendix 9, we further generalize this framework by allowing $W^{i}_{t}$ to follow more general probability distributions $\pi_{t}$ and defining the experimental design as $\Pi=\pi_{0}\times\ldots\times\pi_{T}$. The random variable $W^{i}_{t}$ can take values in either an integer set (e.g., different treatment types) or a continuous set (e.g., varying treatment doses). This generalization distinguishes our work from recent literature on network interference and longitudinal data, which primarily focuses on binary treatments (Arkhangelsky and Imbens, 2023).
For a fixed integer $n^{\text{x}}$, we can also consider covariates (might be observed or not) in the form of a $n^{\text{x}}$ by $N$ matrix $\bm{X}$, where each column (denoted by ${\vec{X}}^{i}:=(X^{1i},\ldots,X^{n^{\text{x}}i})^{\top}$) represents characteristics of unit $i$ (e.g., age and gender). The experimental data thus consists of treatment assignments $\bm{w}{}$, observed outcomes
$\bm{Y}{}(\bm{w}{}):=\bm{Y}{}(\bm{W}{}=\bm{w}{})\,,$
and covariates $\bm{X}$. Therefore, we observe outcomes only under one specific treatment allocation $\bm{w}{}$—a single realization among $2^{N(T+1)}$ distinct potential outcomes, where this number grows exponentially with population size and time horizon. Consequently, estimating causal effects under general interference is impossible due to non-identifiability (Karwa and Airoldi, 2018). To address this challenge, we propose a tractable outcome specification that aligns with and extends the causal message passing framework (Shirani and Bayati, 2024).
@@ 3.2 Potential Outcome Specification
Let $\bm{W}_{t}{}:=[\vec{W}_{0}{}|\ldots|\vec{W}_{t}{}]$ denote the treatment assignments up to time $t$; we represent by $\vec{Y}_{t}{}(\bm{W}_{t}{})=\big{(}Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})\big{)}^{\top}$ the potential outcome vector at time $t$. Consider unknown functions $g_{t}$ and $h_{t}{}$ which operate component-wise, and the expressions $g_{t}\big{(}\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big{)}$ and $h_{t}{}\big{(}\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big{)}$ represent the corresponding column vectors. Given $\vec{Y}_{0}{}(\bm{W}_{0}{})$ as the initial outcome vector, we consider the following specification for potential outcomes:
$\vec{Y}_{t+1}{}(\bm{W}_{t+1}{})=\big{(}{\bf A}+{\bf B}_{t}\big{)}g_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+h_{t}{}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+\vec{\epsilon}^{\;}_{t},\;\;t=0,1,\ldots,T-1,$ (3)
where ${\bf A}$ and ${\bf B}_{t}$ are $N\times N$ unknown matrices capturing the fixed and time-dependent interference effects, respectively. Additionally, $\vec{\epsilon}^{\;}_{t}=\big{(}\epsilon^{1}_{t},\ldots,\epsilon^{N}_{t}\big{)}^{\top}$ is the zero-mean noise term accounting for observation noise. Denoting by ${\text{A}}^{ij}$ and ${\text{B}}^{ij}_{t}$ the element in the $i^{th}$ row and $j^{th}$ column of ${\bf A}$ and ${\bf B}_{t}$, respectively, the value ${\text{A}}^{ij}+{\text{B}}^{ij}_{t}$ quantifies the impact of unit $j$ on unit $i$ at time $t$.
The specification in Eq. (3) captures several aspects of experimental data. It accommodates various types of interference, including treatment spillover effects, carryover effects, peer effects, and autocorrelation (Shirani and Bayati, 2024). Notably, this specification accounts for outcome dynamics by acknowledging the temporal interrelation of units’ outcomes. This contrasts with existing approaches to panel data analysis, which assume that time labels can be shuffled without affecting causal effects (Arkhangelsky and Imbens, 2023).
@@
Next, we analyze the state evolution of the experimental population, characterizing the asymptotic dynamics of unit outcomes. These theoretical foundations establish the basis for developing robust counterfactual estimators in subsequent sections.
@@ 3.3 Experimental State Evolution
In this section, we analyze the distribution of unit outcomes over time. This analysis requires the following two assumptions about the interference matrices. {assumption}[Gaussian interference structure] For all $i,j$, the element ${\text{A}}^{ij}$ in the $i^{th}$ row and $j^{th}$ column of ${\bf A}$ is an independent Gaussian random variable with mean $\mu^{ij}/N$ and variance $\sigma^{2}/N$. Similarly, ${\text{B}}^{ij}_{t}$, the element in the $i^{th}$ row and $j^{th}$ column of ${\bf B}_{t}$, is an independent Gaussian random variable with mean $\mu^{ij}_{t}/N$ and variance $\sigma^{2}_{t}/N$. {assumption}[Convergent interference pattern - informal statement] For all unit $i$ and any time $t$, the elements of vector $(\mu^{i1},\ldots,\mu^{iN})$ admit a weak limit [^6] This means that the empirical distribution of ${\mu^{i1},\ldots,\mu^{iN}}$ converges to a probability distribution as $N$ increases. and, separately, the elements of vector $(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})$ admit a weak limit, where both limits are invariant in $i$. According to Assumptions 3.3 and 3.3, the impact of unit $j$ on unit $i$ is captured by $\mu^{ij}$, $\mu^{ij}_{t}$, and two centered Gaussian random variables. This generalizes the model of Shirani and Bayati, (2024), which assumes i.i.d. interference matrix elements across all units. This extension accommodates more heterogeneous local interactions and varying levels of uncertainty about the interference structure. Specifically, a fully known interference network has exact values for $\mu^{ij}$ and $\mu^{ij}_{t}$ with $\sigma^{2}=0$ and $\sigma^{2}_{t}=0$, while a completely unknown interference means no knowledge of these quantities. Importantly, our estimation method is designed to handle the cases that we have no knowledge of these underlying quantities for implementation.
@@
Consider the case where $\mu^{ij}$ and $\mu^{ij}_{t}$ take values of 00 and $1$, generating time-dependent adjacency matrices of the graph representing the interference structure. If we have high confidence that interactions occur only through these adjacency matrices, we can imagine negligible values for $\sigma^{2}$ and $\sigma^{2}_{t}$. Conversely, significant uncertainty about potential interactions would be reflected in larger values of $\sigma^{2}$ and $\sigma^{2}_{t}$.
@@
The formal version of Assumption 3.3 appears in Assumption 9.4 in Appendix 9.4, where we generalize the condition by allowing greater variation across units. This interference model accommodates complex interaction patterns and admits further generalizations, including extensions to non-Gaussian interference matrices. For detailed discussions, see the Appendix of Shirani and Bayati, (2024).
@@
We then establish the following informal result for the large-sample regime that characterizes how outcome distributions evolve between consecutive time points, with its formal statement presented in Appendix 9.4.
Theorem 3.5 (State evolution - informal statement)
@@
Let $N\to\infty$ and suppose Assumptions 3.3 and 3.3 hold. There exist mappings $f_{t},\;t=0,\ldots,T-1$, such that the distribution of outcomes at time $t+1$ is determined by:
$\mathcal{L}(Y_{t+1}{})=f_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{})\right),$ (4)
@@
where $Y_{t}{}$ denotes the weak limit of unit outcomes $Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})$ at time $t$, $W_{t+1}{}$ follows a Bernoulli distribution with mean $p_{t+1}$, and $\mathcal{L}(\cdot)$ refers to law or distribution of its argument.
@@
The mapping $f_{t}$ in Eq. (4) characterizes how the distribution of outcomes at time t+1 emerges from the previous distribution at time $t$ through the functions $g_{t}$ and $h_{t}{}$, while accounting for both direct treatment effects and indirect effects from the interference matrices and unit covariates. Specifically, in the large-sample limit, even though each unit’s outcome depends on a complex network of interactions, the population-level distribution of outcomes follows a simpler evolution that depends only on the previous distribution and treatment assignment distribution. Building on (Shirani and Bayati, 2024), while extending their results to a broader setting, we refer to the relationship in (4) as the experimental state evolution (SE) equation.
3.4 A First Look at Estimation Through Theory and Practical Constraints
The state evolution equation (4) suggests a natural estimation strategy, motivated by the observation that in many settings, $f_{t}$ remains relatively stable over time. Specifically, we can often decompose $f_{t}$ into a substantial time-invariant component $f$ and a time-varying component.
This decomposition, combined with the fact that empirical distributions from experimental data with many units should approximate the distributions in equation (4), enables a supervised learning approach to estimate $f$. When strong time trends are present, one can first detrend the data to isolate the time-invariant component before applying these learning techniques. Once these mappings are estimated, we can generate any desired counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ through recursive application of Eq. (4) (see Algorithm 3 in Appendix 10). For the sake of building intuition, we will explain a simple version of this, adapted from Shirani and Bayati, (2024).

### v2

@@ 3.1 Experimental Design Framework
We formalize the concept of experimental design by its properties in the large-population limit. Consider the $N\times(T+1)$ treatment matrix $\bm{W}{}$ with entries $W^{i}_{t}$. Each row, $\vec{W^{i}}:=(W^{i}_{0},\dots,W^{i}_{T})$, represents the full history of treatment assignments for unit $i$. We assume that as $N\to\infty$, the empirical distribution of these $N$ row vectors converges weakly to a well-defined probability distribution. We define this limiting distribution, denoted by $\Pi_{T}$, as the experimental design.
In many applications, treatments are binary (e.g., $W^{i}_{t}\in\{0,1\}$), in which case the support of $\Pi_{T}$ is $\{0,1\}^{T+1}$. If the randomization is performed such that the proportion of units treated at time $t$ approaches $p_{t}$, then the marginal expectation of the $t^{th}$ coordinate of a random vector drawn from $\Pi_{T}$ is precisely $p_{t}$. While we use this Bernoulli case for illustration throughout the main text, our framework accommodates more general scenarios, as detailed in Remark 3.1.
@@
Our framework generalizes beyond the binary case. In Appendix 10, we allow for integer (e.g., different treatment types) or continuous-valued (e.g., varying treatment doses) treatments, a setting more general than much of the current literature (see, e.g., the survey by Arkhangelsky and Imbens, (2023)).
Importantly, this definition of the experimental design through its limiting distribution is agnostic to the dependence structure across units or time periods. The assignments $W^{i}_{t}$ and $W^{i^{\prime}}_{t^{\prime}}$ may be arbitrarily correlated for any two distinct units $i,i^{\prime}$ and any time periods $t,t^{\prime}$. This generality encompasses a broad range of experimental designs, including staggered roll-out design (Xiong et al., 2024a ), micro-randomized trials (Li and Wager, 2022a ), and switchback experiments (Hu and Wager, 2022, Bojinov et al., 2023, Jia et al., 2024).
For a fixed integer $n^{\text{x}}$, we can also consider covariates (might be observed or not) in the form of a $n^{\text{x}}$ by $N$ matrix $\bm{X}$, where each column (denoted by ${\vec{X}}^{i}:=(X^{1i},\ldots,X^{n^{\text{x}}i})^{\top}$) represents characteristics of unit $i$ (e.g., age and gender). The experimental data thus consists of treatment assignments $\bm{w}{}$, observed outcomes $\bm{Y}{}(\bm{w}{}):=\bm{Y}{}(\bm{W}{}=\bm{w}{})\,$, and covariates $\bm{X}$. Therefore, we observe outcomes only under one specific treatment allocation $\bm{w}{}$. This represents a single realization among $2^{N(T+1)}$ distinct potential outcomes, where this number grows exponentially with both population size and time horizon. Consequently, estimating causal effects under general interference becomes impossible due to non-identifiability issues (Karwa and Airoldi, 2018). To overcome this challenge, we propose a tractable outcome specification that aligns with and extends the causal message-passing framework (Shirani and Bayati, 2024).
@@ 3.2 Potential Outcome Specification
Let $\bm{W}_{t}{}:=[\vec{W}_{0}{}|\ldots|\vec{W}_{t}{}]$ denote the treatment assignments up to time $t$; we represent by $\vec{Y}_{t}{}(\bm{W}_{t}{})=\big(Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})\big)^{\top}$ the potential outcome vector at time $t$. Consider unknown functions $g_{t}$ and $h_{t}$ which operate component-wise, and the expressions $g_{t}\big(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big)$ and $h_{t}\big(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big)$ represent the corresponding column vectors. Given $\vec{Y}_{0}{}(\bm{W}_{0}{})$ as the initial outcome vector, we consider the following specification for potential outcomes for $t=0,1,\ldots,T-1$:
$\vec{Y}_{t+1}{}(\bm{W}_{t+1}{})=\big({\bf A}+{\bf B}_{t}\big)g_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+h_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+\vec{\epsilon}^{\;}_{t},$ (3)
where $\bf A$} and ${\bf B}_{t}$ are $N\times N$ unknown matrices capturing the fixed and time-dependent interference effects, respectively. Additionally, $\vec{\epsilon}^{\;}_{t}=\big(\epsilon^{1}_{t},\ldots,\epsilon^{N}_{t}\big)^{\top}$ is the zero-mean random variable accounting for observation noise. Denoting by ${\text{A}}^{ij}$ and ${\text{B}}^{ij}_{t}$ the element in the $i^{th}$ row and $j^{th}$ column of $\bf A$} and ${\bf B}_{t}$, respectively, the value ${\text{A}}^{ij}+{\text{B}}^{ij}_{t}$ quantifies the impact of unit $j$ on unit $i$ at time $t$.
The specification in Eq. (3) captures several aspects of experimental data. It accommodates various types of interference, including treatment spillover effects, carryover effects, peer effects, and autocorrelation; Figure 1 of Shirani and Bayati, (2024) illustrates these effects in a simple example. Notably, this specification accounts for outcome dynamics by acknowledging the temporal interrelation of units’ outcomes. This contrasts with existing approaches to panel data analysis, which assume that time labels can be shuffled without affecting causal effects (Arkhangelsky and Imbens, 2023).
@@
Next, we analyze the state evolution of the experimental population, characterizing the asymptotic dynamics of unit outcomes that will provide theoretical foundations for developing robust counterfactual estimators in subsequent sections.
@@ 3.3 Experimental State Evolution
In this section, we analyze the distribution of unit outcomes over time. This analysis requires the following two assumptions about the interference matrices. {assumption}[Gaussian interference structure] For all $i,j$, the element ${\text{A}}^{ij}$ in the $i^{th}$ row and $j^{th}$ column of $\bf A$} is an independent Gaussian random variable with mean $\mu^{ij}/N$ and variance $\sigma^{2}/N$. Similarly, ${\text{B}}^{ij}_{t}$, the element in the $i^{th}$ row and $j^{th}$ column of ${\bf B}_{t}$, is an independent Gaussian random variable with mean $\mu^{ij}_{t}/N$ and variance $\sigma^{2}_{t}/N$. {assumption}[Convergent interference pattern] For all unit $i$ and any time $t$, the elements of vector $(\mu^{i1},\ldots,\mu^{iN})$ admit a weak limit [^5] This means that the empirical distribution of $\{\mu^{i1},\ldots,\mu^{iN}\}$ converges to a probability distribution as $N$ increases. and, separately, the elements of vector $(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})$ admit a weak limit, where both limits are invariant in $i$. We later relax this $i$-invariance in Appendix 10.4-Assumption 10.4. According to Assumptions 3.3 and 3.3, the impact of unit $j$ on unit $i$ is captured by $\mu^{ij}$, $\mu^{ij}_{t}$, and two centered Gaussian random variables. This generalizes the model of Shirani and Bayati, (2024), which assumes i.i.d. interference matrix elements across all units. This extension accommodates more heterogeneous local interactions and varying levels of uncertainty about the interference structure. Specifically, a fully known interference network has exact values for $\mu^{ij}$ and $\mu^{ij}_{t}$ with $\sigma^{2}=0$ and $\sigma^{2}_{t}=0$, while a completely unknown interference means no knowledge of these quantities. Importantly, our estimation method is designed to handle the cases that we have no knowledge of these underlying quantities for implementation.
@@
Consider the case where $\mu^{ij}$ and $\mu^{ij}_{t}$ take values of 0 and $1$, generating time-dependent adjacency matrices of the graph representing the interference structure. If we have high confidence that interactions occur only through these adjacency matrices, we can imagine negligible values for $\sigma^{2}$ and $\sigma^{2}_{t}$. Conversely, significant uncertainty about potential interactions would be reflected in larger values of $\sigma^{2}$ and $\sigma^{2}_{t}$.
@@
The formal version of Assumption 3.3 appears in Appendix 10.4-Assumption 10.4, where we generalize the condition by allowing greater variation across units. Furthermore, while this interference model assumes Gaussian distributions conditional on $\mu^{ij}$ and $\mu^{ij}_{t}$, the appendix of Shirani and Bayati, (2024) discusses how this conditional Gaussianity assumption can be relaxed to accommodate more general distributional forms.
@@
We next present the following informal result for the large-sample regime that characterizes outcome dynamics between consecutive periods, with a formal statement presented in Theorem 4.2.
State evolution - informal statement.
@@
Let $N\to\infty$ and suppose Assumptions 3.3 and 3.3 hold. There exist mappings $f_{t}$, which depend on ${\bf A},{\bf B}_{t},g_{t},h_{t},\bm{X}$ and $\vec{\epsilon}^{\;}_{t}$, $t=0,\ldots,T-1$, such that the distribution of outcomes at time $t+1$ is determined by:
$Y_{t+1}{}\stackrel{{\scriptstyle d}}{{=}}f_{t}\left(Y_{t}{},W_{t+1}{}\right),$ (4)
@@
where $Y_{t}{}$ denotes the weak limit of outcomes $Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})$, $W_{t+1}{}$ follows a Bernoulli distribution with mean $p_{t+1}$, and $\stackrel{{\scriptstyle d}}{{=}}$ refers to equality in distribution.
@@
The mapping $f_{t}$ in Eq. (4) characterizes how aggregate outcomes evolve over time. Specifically, in the large-sample limit, even though each unit’s outcome depends on a complex network of interactions, the population-level distribution of outcomes follows a simpler evolution that depends only on the previous distribution and treatment assignment distribution. Building on Shirani and Bayati, (2024), while extending their results to a broader setting, we refer to the relationship in Eq. (4) as the experimental state evolution (SE) equation.
3.4 General Estimation Theory
The state evolution equation (4) motivates a natural estimation strategy. In large experimental populations, the empirical distributions of observed outcomes at each time period approximate the theoretical distributions specified in Eq. (4). In other words, and loosely speaking, we can obtain approximations for $Y_{t+1}{}$, $Y_{t}{}$, and $W_{t+1}{}$, from experimental data. Leveraging this approximation, we can employ supervised learning methods to estimate the state evolution mappings $f_{t}$. Once estimated, these mappings enable recursive generation of any desired counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ by repeated application of Eq. (4). Our counterfactual estimation problem therefore reduces to consistently estimating $f_{t}$. We present this result below with an informal statement, while providing the rigorous theorem and proof in Theorem 11.1 of Appendix 11.
Theorem 3.5 (Consistency - informal statement)
Suppose that functions $g_{t}$ and $h_{t}$ are affine in their first argument. [^6] There exist functions $g_{t}^{1},g_{t}^{2},h_{t}^{1}$ and $h_{t}^{2}$ such that $g_{t}(y,\cdot)=yg_{t}^{1}(\cdot)+g_{t}^{2}(\cdot)$ as well as $h_{t}(y,\cdot)=yh_{t}^{1}(\cdot)+h_{t}^{2}(\cdot)$. Under certain regularity conditions, given consistent estimation of the mappings $f_{t}$ in the state evolution equation (4), any desired counterfactual evolution can be consistently estimated. Specifically, we can construct an estimator $\widehat{\text{CFE}}_{t}{(\bm{w}{}^{\prime})}$ such that as $N\rightarrow\infty$, for all $t$:
$\widehat{\text{CFE}}_{t}{(\bm{w}{}^{\prime})}-\text{CFE}_{t}{(\bm{w}{}^{\prime})}\xrightarrow{P}0,$
where the convergence is in probability. When the estimation of the mappings $f_{t}$ achieves strong consistency, this convergence can be strengthened to almost sure convergence.
Theorem 3.5 reframes the counterfactual estimation problem as a supervised learning task, and our immediate goal is to estimate the functions $f_{t}$ from observed data. However, we observe only a single experimental scenario with interference-induced dependence; consequently, successful generalization relies on balancing the flexibility of the hypothesis class for functions $f_{t}$ against the available sample size. Inspired by the invariance of treatment propagation mechanisms over time, we pool statistical strength across time and adopt Assumption 3.5. Specifically, this assumption requires that the underlying interaction mechanisms remain stable over the experimentation horizon. In other words, we can decompose $f_{t}$ into a substantial time-invariant component that captures the core causal mechanisms and a smaller time-varying component that accounts for temporal variations. Nevertheless, as we discuss in Section 3.5, Assumption 3.5 alone may not suffice under data scarcity; Sections 4 and 5 propose remedies. {assumption}[Weakly additive time trend] Considering the state evolution equation (4), there exist mappings $f$ and $\tilde{f}_{t}$ such that for each $t$:
$f_{t}\left(Y_{t}{},W_{t+1}{}\right)=f\left(Y_{t}{},W_{t+1}{}\right)+\tilde{f}_{t}\left(Y_{t}{}\right)$ (5)
Under Assumption 3.5, our primary objective becomes estimating the time-invariant mapping $f$, which captures the essential treatment effects according to Eq. (5). When temporal variations are minimal (e.g., when $\tilde{f}_{t}$ is a random function and $\mathbb{E}[\tilde{f}_{t}\left(Y_{t}{}\right)]=0$), we can safely ignore the time-dependent component $\tilde{f}_{t}$ and directly estimate $f$ from the available data (Figure 3). To build intuition, we present a simplified version of this approach below, adapted from Shirani and Bayati, (2024). Conversely, when strong temporal patterns are present, we first estimate and remove these trends by modeling $\tilde{f}_{t}$, thereby isolating the time-invariant component $f$. We can then apply supervised learning techniques to estimate this core mapping. We defer the technical details of this more complex case to Appendix 12.
Figure 3: Estimation strategy: Experimental data is used to estimate state evolution mappings through supervised learning, which are then applied recursively to generate desired counterfactual evolutions. Treatment allocations $\bm{w}$ and $\bm{w^{\prime}}$ share identical initial columns, serving as initialization for our recursive approach. While observations are collected at the unit level, the estimated counterfactuals represent population-level quantities.